Financial analysts forecast Crestwood Equity Partners (CEQP) growth for the future to be 3.1 percent. The firm just paid a $1.22 dividend. What is the value of their stock when the required rate of return is 13.9 percent?
Answer #1

Ans $ 11.65

P0 = Price of Share
D1 = Current Dividend
Ke = Cost of Equity
g = growth rate
P0 = D1 / (Ke - g)
P0 = 1.2578 / (13.9%- 3.1%)
P0 = 11.65
D1 = D0* (1 + g)
D1 = 1.22* (1 + 3.1%)
D1 = 1.2578